40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 11346     Entry time: Fri Jun 5 11:59:59 2015
Author: ericq 
Type: Bureaucracy 
Category: General 
Subject: Maintenance Tasks, IFO upgrades 

At wednesday's meeting, Rana, Koji, Steve, and I started making a list of maintenence tasks that should be done/checked on a regular basis. The actual scheduling of these has not yet been considered. They include:

  • N2 Tank pressure / cylinder replacement
  • Headlamp, walkie talkie battery recharge
  • Workstation software updates
  • Coffee bean and filters
  • Multimeter battery levels
  • Sorensen DC power supply voltage settings and current draws
  • UPS' status (Vacuum, NFS host, workstations)
  • SR560s, battery powered scopes plugged in
  • Rack Fuses intact
  • Take pictures of electronics racks, optical tables
  • Replace PSL HEPA filter

Next, we brainstormed work that can be done to improve the interferometer performance, and what order/precedence they should take. 

In the end, it was decided that the plan for the next few weeks was to focus on improving the ALS noise levels, and, more importantly, seeking to make the performance more consistent. We need to know what is limiting us, and what we extent we can expect to improve things. To this end, I am working on reviving a ALS noise budget; using the noise budget from the green locking paper to inform a simulinkNB kind of thing.

Here are all of the items we listed during this brainstorming session.

Some near-term/priority tasks are:

  • Installing the accelerometers near MC1 and 2
  • Installing green steering PZT mirrors at the Y end table, commission dither alignment
  • Improving the X end green mode matching
  • ALS noise budgeting
  • Upgrade the realtime system to RCG v2.9

More down the line, other things we thought about were:

  • Cleanup of bench power supplies (FSS box has one, where else?)
  • Fixing the ETMX suspension issues 
  • Upgrade SOS suspension code to the appropriate aLIGO block
  • Upgrade to the green PDH electronics
  • Understanding / tuning the FSS servo laser PZT vs. PC crossover
  • Undestanding / tuning the 11MHz vs 55MHz modulation phase
  • Replacing the slow vxworks machines with the Acromag setup Aiden has set up for the CTN lab
  • QPD upgrade
  • New/better green beatbox
  • Finalize the manifestation of the IR beat control (Freq counter vs. fast DFD)
  • Explore the idea of using an analog output of the ALS beatbox as fast input to the CM board
  • Replace triple resonant EOM driving circuit with double resonant one
  • X end table layout/enclosure upgrade
ELOG V3.1.3-